Application 

1. Hydrogen gas can easily be generated using the reaction of magnesium with hydrochloride acid HCL,for a reaction, 500ml of hydrogen gas was collected by water displacement. Calculate the mass of magnesium required to produce "dry" hydrogen gas a total pressure of 125 kPa and 22°C? 

2. Large quantities of methane gas are being released from arctic waters as water temperature rise due to global warming. A methane bubble has a volume of 2000L at 10 atm of pressure and 4.0°C. What will the volume of the bubble be at 10°C and 102 kPa

3. Carbon dioxide in high levels can cause death. One method of removing carbon dioxide from the air of enclosed compartments involves this reaction 

2NaOH + CO2 (g) = Na2CO (s) + H2O (l)

What mass of sodium hydroxide is required to remove all the carbon dioxide in the air over a 5 hr period, if the volume of CO2 produced is 1.25L/hr at a temperature of 22°C And a pressure of 101.8kPa?


1
Expert's answer
2021-05-12T06:29:16-0400

1.

"Mg+ 2HCl \to MgCl_2+ H_2"

"PV = nRT\\
125000×500×10^{-6}= n×8.314×(273.15+22)"

"62.5 = 2453.88n"


"n = \dfrac{62.5}{2453.88}= 0.0255\ moles"


mass of hydrogen = 0.0255 mol × 2g/mol = 0.051g


2.

"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"


"\dfrac{10×102×2000}{277}=\dfrac{102×V_2}{283}"


"V_2 =\dfrac{283×10×2000}{277}=20433.21\ L"


3.

"2NaOH + CO_2 \to Na_2CO + H_2O"


Volume of CO2 produced over 5 hours is 6.25L

"PV = nRT\\
101800× 6.25× 10^{-3} = n×8.314× (273+22)\\
636.25 = 2452.63n"


"n = \dfrac{636.25}{2452.63}= 0.26\ moles"


According to the reaction above,

0.26 moles of CO2 requires 0.52 moles of NaOH to go into completion.


mass of NaOH required = 0.52 moles × 40 g/mol = 20.8g